# Neural Post-Einsteinian Framework for Efficient Theory-Agnostic Tests of General Relativity with Gravitational Waves

### v1

@@ Abstract
The parametrized post-Einsteinian (ppE) framework and its variants are widely used to probe gravity through gravitational-wave tests that apply to a large class of theories beyond general relativity. However, the ppE framework is not truly theory-agnostic as it only captures certain types of deviations from general relativity: those that admit a post-Newtonian series representation in the inspiral of coalescencing compact objects. Moreover, each type of deviation in the ppE framework has to be tested separately, making the whole process computationally inefficient and expensive, possibly obscuring the theoretical interpretation of potential deviations that could be detected in the future. We here present the neural post-Einsteinian (npE) framework, an extension of the ppE formalism that overcomes the above weaknesses using deep-learning neural networks. The core of the npE framework is a variantional autoencoder that maps the discrete ppE theories into a continuous latent space in a well-organized manner. This design enables the npE framework to test many theories simultaneously and to select the theory that best describes the observation in a single parameter estimation run. The smooth extension of the ppE parametrization also allows for more general types of deviations to be searched for with the npE model. We showcase the application of the new npE framework to future tests of general relativity with the fifth observing run of the LIGO-Virgo-KAGRA collaboration. In particular, the npE framework is demonstrated to efficiently explore modifications to general relativity beyond what can be mapped by the ppE framework, including modifications coming from higher-order curvature corrections to the Einstein-Hilbert action at high post-Newtonian order, and dark-photon interactions in possibly hidden sectors of matter that do not admit a post-Newtonian representation.
@@ I Introduction
Tests of GR with GW data generally fall into two categories: consistency tests and parametrized tests [28]. Consistency tests, such as the residual test and the inspiral-merger-ringdown consistency test, search for possible violations of GR by asking how consistent the observed signal is with the prediction of GR. Parametrized tests, on the other hand, search for possible non-GR effects that are encoded by certain parameters in the gravitational waveform model. Although these parameters are related to specific non-GR theories, parameterized tests can still be theory-agnostic, as long as the parametrization covers the predictions of many non-GR models. One successful approach is provided by the parametrized post-Einstein (ppE) framework [29, 30, 31, 32], which introduces small deformations to the GR waveform that are motivated by generic, theoretical considerations and by a broad class of non-GR theories.
@@
The ppE framework is the general framework that the LVK’s “parametrized inspiral tests of GR” is built from [33, 34, 35, 36, 10, 37, 38, 39, 40, 41, 28, 42]. In these tests, the deformation is given by the leading-order modification to the post-Newtonian (PN) expansion111 This is an expansion of the inspiral waveform in powers of the binary velocity, which can be reexpressed in terms of the GW frequency through the PN version of Kepler’s third law [43]. of the phase of the GR waveform. Such a prescription introduces two parameters: a ppE exponent or ppE index representing the PN order at which the modification first takes place, and a ppE coefficient representing the magnitude of the deformation at that given PN order. In terms of a physical interpretation, the ppE index maps to the type of non-GR modification one wishes to consider, and the ppE coefficient maps to the strength of the modification, including the coupling constants of modified gravity theories [29, 13, 44, 45]. To run ppE or parameterized inspiral tests on GW data, one fixes the ppE index at a chosen value and estimates the ppE coefficient via Bayesian inference. The test is then made theory-agnostic by repeating the estimation for all ppE indices consistent with a non-GR theoretical prediction [10, 38, 39, 40].
@@
Despite being standard, the ppE or parametrized inspiral test is unsatisfying for two reasons. First, the ppE test is not truly theory-agnostic. The ppE deformation applied to an inspiral waveform only captures modification at leading PN order. This means the modification must admit a PN expansion in the first place. Counterexamples exist, like the GWs generated by a binary in massive scalar-tensor theories [46] or with dark-photon interactions [47], in which non-GR modifications activate suddenly, and thus, cannot be represented as a simple power of velocity. Even for modifications that are PN-expandable, a modeling bias can exist when certain higher-PN-order effects are omitted. Although one may formally add higher PN order terms to the ppE deformation, they must be accompanied with more coefficient parameters. Bayesian parameter estimation with this expanded parameter space will lead to uninformative results unless a certain marginalization procedure is implemented [48]. The second reason ppE inspiral tests are not ideal is their high computational expense, because of the need for repeating parameter estimation for many ppE indices. This issue may be mitigated by accelerating each estimation run with a hybridized sampling technique [49], but still, one must repeat the analysis for many ppE indices. A trans-dimensional reversible-jump Markov-chain Monte-Carlo algorithm can alleviate the latter [50, 51], but at an increased computational cost. We also note that a principal component analysis allows multiple PN terms to be estimated at once to search for potential deviations from GR [52, 53, 54], but the deviations constrained or found this way cannot be mapped back to any specific non-GR theory, and hence, imply nothing about fundamental physics.
@@ II.1 CBC Process and GW Parameter Estimation
Discussion in later sections will focus on quasi-circular and non-precessing, binary black holes (BBH), and will use the IMRPhenomD waveform [61, 62] to describe GWs from such sources in GR. The IMRPhenomD waveform models the GW response function in the frequency domain via
@@
where $p(\{\Theta_{\alpha}\}|d)$ is the posterior for the set of CBC parameters $\Theta_{\alpha}$ given the the GW data $d$, $p(\{\Theta_{\alpha}\})$ is the prior of the parameters, and $p(d|\{\Theta_{\alpha}\})$ is the likelihood function. Assuming that the detector noise is additive and Gaussian, the likelihood can be further expressed as:
@@ II.2 GW Tests of GR with the ppE Framework
We note that the test described above is theory-specific. A more theory-agnostic test can be built from the ppE framework, like those performed by the LVK [10, 38, 39, 40, 41, 28]. In this scheme, the GR waveform is deformed by terms that are motivated by generic theoretical considerations and cover a broad class of non-GR theories. Let us then follow the prescription of [13] and consider only modifications to the frequency-domain phase of the inspiral GW. The ppE waveform can then be writen as
@@
To run a ppE or parameterized test of GR, one chooses a list of integer $b_{\mathrm{ppE}}$’s and estimate $\beta_{\mathrm{ppE}}$ for each one of them. If any of the $\beta_{\mathrm{ppE}}$ posteriors is statistically incompatible with zero, a deviation from GR can be claimed under the physical mechanisms implied by the corresponding $b_{\mathrm{ppE}}$. If GR is not rejected by the posterior, the constraint on each $\beta_{\mathrm{ppE}}$ can be mapped to constraints on the theories associated. The ppE framework is relatively efficient, because a large number of non-GR theories are “compressed” into a smaller (and finite) set of ppE indices. Moreover, if new theories are developed in the future, chances are that they can be represented in the ppE way at leading PN order, and that constraints on those ppE indices may already have been obtained.
@@
Despite having been widely adopted, [33, 34, 35, 36, 10, 37, 38, 39, 40, 41, 28, 42] the ppE framework can be unsatisfactory for two main reasons. First, ppE tests are not truly theory-agnostic. Equation (10) only makes sense for theories that admit a PN expansion for their description of the inspiral. Counterexamples exist, like the massive scalar-tensor theories [46] and theories with dark-photon interactions in the hidden sectors of matter [47], in which this is not the case. In these theories, the gravitational dynamics is affected by the competition between GW quadrupole emission, occurring at the binary scale, and scalar emission, which can be dynamically activated at a certain (Compton wavelength) scale associated with the mass of the scalar field. A typical consequence of this is that the phase of the inspiral waveform acquires modifications that are proportional to a Heaviside step function in frequency (to be discussed in more detail in Sec. IV.5). This step function introduces sudden modifications to the waveform phase, which cannot be described as a Taylor series in velocity. Moreover, even for theories that admit PN-expandable modifications, a modeling bias can exist when certain higher-PN-order effects are omitted. Although one may formally add higher PN order terms to Eq. (10), they must be accompanied with more coefficient parameters, and the prior range of each coefficient must be carefully regularized in order for Bayesian parameter estimation to produce informative results after marginalizing over the higher-PN-order terms [48].
@@
The second reason ppE inspiral tests are not ideal is their high computational expense. Although the ppE framework is more efficient than separately testing each specific theory available in the literature, one still has to run Bayesian parameter estimation for every element in a list of ppE indices, repeating parameter estimation many times. This issue may be mitigated by a sampling technique that hybridizes one expensive estimation in GR with a set of cheaper estimations that include ppE deviations [49], but still one must repeat the analysis for many ppE indices. Though the latter can be alleviated with a trans-dimensional reversible-jump Markov-chain Monte Carlo algorithm that enables sampling across parameter spaces with varying dimensions [50, 51] (such as transitions between GR models and non-GR ppE models), such an algorithm is complex and may elevate computational expenses.
@@ III.2 A Sketch of the npE Framework
Because we intend to use the npE framework for tests of GR through Bayesian inference, we expect the vector $z_{i}$ in $\Phi_{\mathrm{npE}}$ to be accompanied by a simple prior that describes a uniform distribution of theories for modifications of different magnitude. An example of such a simple prior, and the choice we will make in this paper, is that of a flat distribution bounded by the unit hypersphere centered at $z_{i}=0$. We choose $z_{i}=0$ to represent GR, and $\lVert z_{i}\rVert=1$ to represent a “prior boundary,” associated with the largest modifications of interest. In this paper, we will require that the modifications to GR be small deformations (as in the ppE framework), and thus, we will not allow these corrections to be larger than the GR terms in the GW phase themselves; this EFT condition sets the size of the prior, which we shall refer to as the EFT prior boundary. With this in mind, the magnitude $\lVert z_{i}\rVert$ represents the size of the modification, $\Phi_{\mathrm{npE}}\propto\lVert z_{i}\rVert$, while the direction in the latent space $n_{i}=z_{i}/\lVert z_{i}\rVert$ represents the type of GR modification.
@@
In our npE phase model, the shape function will be learned through a VAE, which we will train first. The scale function will be learned through a separate neural network, which we refer to as the secondary network and which we will train later, treating the latent parametrization as already completed by the trained VAE. Such a separation of the training processes is beneficial when the prior boundary is chosen to reflect the most stringent constraints on theories of interest. If these constraints are tightened by future observations, one only needs to re-train the secondary network to update the scale function, while the VAE for the shape function does not necessarily require an update.
@@
Figure 2: Ways to connect the npE networks to (a) model the npE phase modification $\Phi_{\mathrm{npE}}$ with the latent parameters $z_{i}$ and the source parameters $\Xi_{\alpha}$, (b) train the VAE for the shape function $S$ given a data set of non-GR phase modifications $\Phi_{\mathrm{mod}}$, and (c) train the secondary network for the scale function $T$ given $\Phi_{\mathrm{mod}}$ and $\Xi_{\alpha}$ in the same data set. The quantity $n_{i}$ represents a unit vector in the latent space, and $\mu_{i}$ and $\sigma_{i}$ are the mean and standard deviation of the encoder output, respectively. Each rounded rectangle represents an operation, either predefined (uncolored) or provided by a neural network (colored). For the predefined operations, $\mathcal{N}$ represents a random draw from a normal distribution, and $\frac{(\cdot)}{\lVert\cdot\rVert}$ means vector normalization. For the neural networks, the name of the network is italic if the network parameters are being fitted, and is roman if the network is treated as a known function and no parameter update takes place.
@@ III.3 Modified Gravity Data Set for the npE Networks
$\beta_{\mathrm{ppE,max}}=\left\{\begin{aligned} &|\phi_{b+5}|\,\eta^{-(b+5)/5},&\,&b>=-5,\\ &|\phi_{0}|\,(\pi\mathcal{M}f_{\mathrm{low}})^{-(b+5)/3},&\,&b<-5,\end{aligned}\right.$ (17)
@@
To numerically represent $\Phi_{\mathrm{mod}}^{(a)}(f)$, we choose a grid of $f$ and evaluate $\Phi_{\mathrm{mod}}^{(a)}$ as a vector composed of phases at each grid point, i.e.
@@ III.4 The VAE Network for the Shape Function
We connect the encoder and the decoder in the latent space, and the output of the decoder to the training data as follows. First, we require $\mu_{i}^{(a)}$ to be a unit vector, i.e. $\lVert\mu_{i}^{(a)}\rVert=1$. Second, we require that the random sample, taken from distributions associated with $\mu_{i}^{(a)}$ and $\sigma_{i}^{(a)}$, be generated from
@@
The encoder and the decoder will be neural networks, but before we define them, let us first discuss some npE specific choices for the network designs. First, because $\Phi_{\mathrm{ppE}}$ is modelled by two non-GR parameters $(b_{\mathrm{ppE}},\beta_{\mathrm{ppE}})$, we choose the VAE latent space to also be two-dimensional, which we find to be sufficient to faithfully represent these ppE theories. Based on that, we also require the two standard deviations of the encoder to be degenerate, i.e. $\sigma_{1}=\sigma_{2}$. This logically makes sense because the true VAE parametrization only takes place on the unit circle, which is just 1-dimensional.
@@
where $U_{n}$ and $V_{n}$ are to be implemented as subnetworks of the decoder. This choice makes our shape network physically-informed, allowing it to emulate any power-law behavior in the training set and recover the standard ppE modifications, while a generic network cannot. Equation (24) also formally addresses how the decoder output is interpolated in the frequency grid. The number of pseudo-PN terms $N_{p}$ is a hyperparameter of the VAE. When $N_{p}=1$, Eq. (24) resembles a naïve extension of the ppE framework, i.e. upgrading the $b_{\mathrm{ppE}}$ from an integer to a real number. When $N_{p}\geq 2$, Eq. (24) resembles an advanced ppE framework with higher-PN corrections included. In this case, we expect the VAE to learn a latent space with fewer dimensions as compared to the total number of PN coefficients to be otherwise included, and distribute a different mixture of PN orders with appropriate prior weights. In this paper, we will fix $N_{p}=2$, but other choices can be studied in the future. Note that, usually, the decoder and the encoder are designed to mirror each other, but here we do not require this property. In particular, we do not ask the encoder to know anything about the pseudo-PN expansion of the decoder, because we wish the encoder to be able to handle more generic modifications.
Let us now specify the lower-level implementation of the VAE. We introduce three neural mappings, ${\mathcal{E}}\!\!:\,\mathbb{R}^{640}\mapsto\mathbb{R}^{3}$, $\mathcal{D}_{U}\!\!:\,\mathbb{R}^{2}\mapsto\mathbb{R}^{2}$ and $\mathcal{D}_{V}\!\!:\,\mathbb{R}^{2}\mapsto\mathbb{R}^{2}$. The neural map ${\mathcal{E}}$ takes in 640 real numbers associated with the modified phase at each frequency grid point, and it returns 3 real numbers associated with the 2 means of the latent space and the 1 standard deviation. The neural map $\mathcal{D}_{U}$ and $\mathcal{D}_{V}$, then, take 2 real numbers in the 2-D latent space and outputs two real numbers to be used as amplitudes and exponents in the pseudo-PN representation of the shape function. Each of these neural maps contains 4 fully connected hidden layers, and each hidden layer contains 512 neurons taking the ReLU activation. Therefore, the entire encoder-decoder network contains 3,486,727 parameters that we must train by minimizing the loss function.

### v2

@@ Abstract
The parametrized post-Einsteinian (ppE) framework and its variants are widely used to probe gravity through gravitational-wave tests that apply to a large class of theories beyond general relativity. However, the ppE framework is not truly theory-agnostic as it only captures certain types of deviations from general relativity: those that admit a post-Newtonian series representation in the inspiral of coalescing compact objects. Moreover, each type of deviation in the ppE framework has to be tested separately, making the whole process computationally inefficient and expensive, possibly obscuring the theoretical interpretation of potential deviations that could be detected in the future. We here present the neural post-Einsteinian (npE) framework, an extension of the ppE formalism that overcomes the above weaknesses using deep-learning neural networks. The core of the npE framework is a variational autoencoder that maps the discrete ppE theories into a continuous latent space in a well-organized manner. This design enables the npE framework to test many theories simultaneously and to select the theory that best describes the observation in a single parameter estimation run. The smooth extension of the ppE parametrization also allows for more general types of deviations to be searched for with the npE model. We showcase the application of the new npE framework to future tests of general relativity with the fifth observing run of the LIGO-Virgo-KAGRA collaboration. In particular, the npE framework is demonstrated to efficiently explore modifications to general relativity beyond what can be mapped by the ppE framework, including modifications coming from higher-order curvature corrections to the Einstein-Hilbert action at high post-Newtonian order, and dark-photon interactions in possibly hidden sectors of matter that do not admit a post-Newtonian representation.
@@ I Introduction
Tests of GR with GW data generally fall into two categories: consistency tests and parametrized tests [28]. Consistency tests, such as the residual test and the inspiral-merger-ringdown consistency test, search for possible violations of GR by asking how consistent the observed signal is with the prediction of GR. Parametrized tests, on the other hand, search for possible non-GR effects that are encoded by certain parameters in the gravitational waveform model. Although these parameters are related to specific non-GR theories, parametrized tests can still be theory-agnostic, as long as the parametrization covers the predictions of many non-GR models. One successful approach is provided by the parametrized post-Einstein (ppE) framework [29, 30, 31, 32], which introduces small deformations to the GR waveform that are motivated by generic, theoretical considerations and by a broad class of non-GR theories.
@@
The ppE framework is the general framework that the LVK’s “parametrized inspiral tests of GR” is built from [33, 34, 35, 36, 10, 37, 38, 39, 40, 41, 28, 42]. In these tests, the deformation is given by the leading-order modification to the post-Newtonian (PN) expansion111 This is an expansion of the inspiral waveform in powers of the binary velocity, which can be reexpressed in terms of the GW frequency through the PN version of Kepler’s third law [43]. of the phase of the GR waveform. Such a prescription introduces two parameters: a ppE exponent or ppE index representing the PN order at which the modification first takes place, and a ppE coefficient representing the magnitude of the deformation at that given PN order. In terms of a physical interpretation, the ppE index maps to the type of non-GR modification one wishes to consider, and the ppE coefficient maps to the strength of the modification, including the coupling constants of modified gravity theories [29, 13, 44, 45]. To run ppE or parametrized inspiral tests on GW data, one fixes the ppE index at a chosen value and estimates the ppE coefficient via Bayesian inference. The test is then made theory-agnostic by repeating the estimation for all ppE indices consistent with a non-GR theoretical prediction [10, 38, 39, 40].
@@
Despite being standard, the ppE or parametrized inspiral test is unsatisfying for two reasons. First, the ppE test is not truly theory-agnostic. The ppE deformation applied to an inspiral waveform only captures modification at leading PN order. This means the modification must admit a PN expansion in the first place. Counterexamples exist, like the GWs generated by a binary in massive scalar-tensor theories [46, 47, 48] or with dark-photon interactions [49], in which non-GR modifications activate suddenly, and thus, cannot be represented as a simple power of velocity. Even for modifications that are PN-expandable, a modeling bias can exist when certain higher-PN-order effects are omitted. Although one may formally add higher PN order terms to the ppE deformation, they must be accompanied with more coefficient parameters. Bayesian parameter estimation with this expanded parameter space will lead to uninformative results unless a certain marginalization procedure is implemented [50]. The second reason ppE inspiral tests are not ideal is their high computational expense, because of the need for repeating parameter estimation for many ppE indices. This issue may be mitigated by accelerating each estimation run with a hybridized sampling technique [51], but still, one must repeat the analysis for many ppE indices. A transdimensional reversible-jump Markov-chain Monte-Carlo algorithm can alleviate the latter [52, 32], but at an increased computational cost. We also note that a principal component analysis allows multiple PN terms to be estimated at once to search for potential deviations from GR [53, 54, 55], but the deviations constrained or found this way cannot be mapped back to any specific non-GR theory, and hence, imply nothing about fundamental physics.
@@ II.1 CBC process and GW parameter estimation
Discussion in later sections will focus on quasicircular and nonprecessing, binary black holes (BBH), and will use the IMRPhenomD waveform [62, 63] to describe GWs from such sources in GR. The IMRPhenomD waveform models the GW response function in the frequency domain via
@@
where $p(\{\Theta_{\alpha}\}|d)$ is the posterior for the set of CBC parameters $\Theta_{\alpha}$ given the GW data $d$, $p(\{\Theta_{\alpha}\})$ is the prior of the parameters, and $p(d|\{\Theta_{\alpha}\})$ is the likelihood function. Assuming that the detector noise is additive and Gaussian, the likelihood can be further expressed as:
@@ II.2 GW tests of GR with the ppE Framework
We note that the test described above is theory-specific. A more theory-agnostic test can be built from the ppE framework, like those performed by the LVK [10, 38, 39, 40, 41, 28]. In this scheme, the GR waveform is deformed by terms that are motivated by generic theoretical considerations and cover a broad class of non-GR theories. Let us then follow the prescription of [13] and consider only modifications to the frequency-domain phase of the inspiral GW. The ppE waveform can then be written as
@@
To run a ppE or parametrized test of GR, one chooses a list of integer $b_{\mathrm{ppE}}$’s and estimate $\beta_{\mathrm{ppE}}$ for each one of them. If any of the $\beta_{\mathrm{ppE}}$ posteriors is statistically incompatible with zero, a deviation from GR can be claimed under the physical mechanisms implied by the corresponding $b_{\mathrm{ppE}}$. If GR is not rejected by the posterior, the constraint on each $\beta_{\mathrm{ppE}}$ can be mapped to constraints on the theories associated. The ppE framework is relatively efficient, because a large number of non-GR theories are “compressed” into a smaller (and finite) set of ppE indices. Moreover, if new theories are developed in the future, chances are that they can be represented in the ppE way at leading PN order, and that constraints on those ppE indices may already have been obtained.
@@
Despite having been widely adopted, [33, 34, 35, 36, 10, 37, 38, 39, 40, 41, 28, 42] the ppE framework can be unsatisfactory for two main reasons. First, ppE tests are not truly theory-agnostic. Equation (10) only makes sense for theories that admit a PN expansion for their description of the inspiral. Counterexamples exist, like the massive scalar-tensor theories [46, 47, 48] and theories with dark-photon interactions in the hidden sectors of matter [49], in which this is not the case. In these theories, the gravitational dynamics is affected by the competition between GW quadrupole emission, occurring at the binary scale, and scalar emission, which can be dynamically activated at a certain (Compton wavelength) scale associated with the mass of the scalar field. A typical consequence of this is that the phase of the inspiral waveform acquires modifications that are proportional to a Heaviside step function in frequency (to be discussed in more detail in Sec. IV.5). This step function introduces sudden modifications to the waveform phase, which cannot be described as a Taylor series in velocity. Moreover, even for theories that admit PN-expandable modifications, a modeling bias can exist when certain higher-PN-order effects are omitted. Although one may formally add higher PN order terms to Eq. (10), they must be accompanied with more coefficient parameters, and the prior range of each coefficient must be carefully regularized in order for Bayesian parameter estimation to produce informative results after marginalizing over the higher-PN-order terms [50].
@@
The second reason ppE inspiral tests are not ideal is their high computational expense. Although the ppE framework is more efficient than separately testing each specific theory available in the literature, one still has to run Bayesian parameter estimation for every element in a list of ppE indices, repeating parameter estimation many times. This issue may be mitigated by a sampling technique that hybridizes one expensive estimation in GR with a set of cheaper estimations that include ppE deviations [51], but still one must repeat the analysis for many ppE indices. Though the latter can be alleviated with a transdimensional reversible-jump Markov-chain Monte Carlo algorithm that enables sampling across parameter spaces with varying dimensions [52, 32] (such as transitions between GR models and non-GR ppE models), such an algorithm is complex and may elevate computational expenses.
@@ III.2 A sketch of the npE framework
Because we intend to use the npE framework for tests of GR through Bayesian inference, we expect the vector $z_{i}$ in $\Phi_{\mathrm{npE}}$ to be accompanied by a simple prior that describes a uniform distribution of theories for modifications of different magnitude. An example of such a simple prior, and the choice we will make in this paper, is that of a flat distribution bounded by the unit hypersphere centered at $z_{i}=0$. We choose $z_{i}=0$ to represent GR, and $\lVert z_{i}\rVert=1$ to represent a “prior boundary,” associated with the largest modifications of interest. In this paper, we will require that the modifications to GR be small deformations (as in the ppE framework), and thus, we will not allow these corrections to be larger than the GR terms in the GW phase themselves; this EFT condition sets the size of the prior, which we shall refer to as the EFT prior boundary [^2] Note of caution is now necessary. By “EFT boundary”, we do not here mean a cut-off boundary beyond which the energies probed are beyond the regime of validity of the effective theory. Rather, we mean a boundary at which the GR modifications cease to be perturbatively small on average over all source parameters.. With this in mind, the magnitude $\lVert z_{i}\rVert$ represents the size of the modification, $\Phi_{\mathrm{npE}}\propto\lVert z_{i}\rVert$, while the direction in the latent space $n_{i}=z_{i}/\lVert z_{i}\rVert$ represents the type of GR modification.
@@
In our npE phase model, the shape function will be learned through a VAE, which we will train first. The scale function will be learned through a separate neural network, which we refer to as the secondary network and which we will train later, treating the latent parametrization as already completed by the trained VAE. Such a separation of the training processes is beneficial when the prior boundary is chosen to reflect the most stringent constraints on theories of interest. If these constraints are tightened by future observations, one only needs to retrain the secondary network to update the scale function, while the VAE for the shape function does not necessarily require an update.
@@
Figure 2: Ways to connect the npE networks to (a) model the npE phase modification $\Phi_{\mathrm{npE}}$ with the latent parameters $z_{i}$ and the source parameters $\Xi_{\alpha}$, (b) train the VAE for the shape function $S$ given a dataset of non-GR phase modifications $\Phi_{\mathrm{mod}}^{\mathrm{train}}$, and (c) train the secondary network for the scale function $T$ given $\Phi_{\mathrm{mod}}^{\mathrm{train}}$ and $\Xi_{\alpha}^{\mathrm{train}}$ in the same dataset. The quantity $n_{i}$ represents a unit vector in the latent space, and $\mu_{i}$ and $\sigma_{i}$ are the mean and standard deviation of the encoder output, respectively. Each rounded rectangle represents an operation, either predefined (uncolored) or provided by a neural network (colored). For the predefined operations, $\mathcal{N}$ represents a random draw from a normal distribution, and $\frac{(\cdot)}{\lVert\cdot\rVert}$ means vector normalization. For the neural networks, the name of the network is italic if the network parameters are being fitted, and is roman if the network is treated as a known function and no parameter update takes place.
@@ III.3 Modified gravity dataset for the npE networks
$\beta_{\mathrm{ppE,max}}=\left\{\begin{aligned} &|\phi_{b+5}|\,\eta^{-(b+5)/5},&\,&b\geq-5,\\ &|\phi_{0}|\,(\pi\mathcal{M}f_{\mathrm{low}})^{-(b+5)/3},&\,&b<-5,\end{aligned}\right.$ (17)
@@
To numerically represent $\Phi_{\mathrm{mod}}^{(a)}(f)$, we choose a grid of $f$ and evaluate $\Phi_{\mathrm{mod}}^{(a)}$ as a vector composed of phases at each grid point, i.e.,
@@ III.4 The VAE network for the shape function
We connect the encoder and the decoder in the latent space, and the output of the decoder to the training data as follows. First, we require $\mu_{i}^{(a)}$ to be a unit vector, i.e., $\lVert\mu_{i}^{(a)}\rVert=1$. Second, we require that the random sample, taken from distributions associated with $\mu_{i}^{(a)}$ and $\sigma_{i}^{(a)}$, be generated from
@@
The encoder and the decoder will be neural networks, but before we define them, let us first discuss some npE specific choices for the network designs. First, because $\Phi_{\mathrm{ppE}}$ is modeled by two non-GR parameters $(b_{\mathrm{ppE}},\beta_{\mathrm{ppE}})$, we choose the VAE latent space to also be two-dimensional, which we find to be sufficient to faithfully represent these ppE theories. Based on that, we also require the two standard deviations of the encoder to be degenerate, i.e. $\sigma_{1}=\sigma_{2}$. This logically makes sense because the true VAE parametrization only takes place on the unit circle, which is just 1-dimensional.
@@
where $U_{n}$ and $V_{n}$ are to be implemented as subnetworks of the decoder. This choice makes our shape network physically informed, allowing it to emulate any power-law behavior in the training set and recover the standard ppE modifications, while a generic network cannot. Equation (24) also formally addresses how the decoder output is interpolated in the frequency grid. The number of pseudo-PN terms $N_{p}$ is a hyperparameter of the VAE. When $N_{p}=1$, Eq. (24) resembles a naïve extension of the ppE framework, i.e. upgrading the $b_{\mathrm{ppE}}$ from an integer to a real number. When $N_{p}\geq 2$, Eq. (24) resembles an advanced ppE framework with higher-PN corrections included. In this case, we expect the VAE to learn a latent space with fewer dimensions as compared to the total number of PN coefficients to be otherwise included, and distribute a different mixture of PN orders with appropriate prior weights. In this paper, we will fix $N_{p}=2$, but other choices can be studied in the future. Note that, usually, the decoder and the encoder are designed to mirror each other, but here we do not require this property. In particular, we do not ask the encoder to know anything about the pseudo-PN expansion of the decoder, because we wish the encoder to be able to handle more generic modifications.
Let us now specify the lower-level implementation of the VAE. We introduce three neural mappings, ${\mathcal{E}}\!\!:\,\mathbb{R}^{640}\mapsto\mathbb{R}^{3}$, $\mathcal{D}_{U}\!\!:\,\mathbb{R}^{2}\mapsto\mathbb{R}^{2}$ and $\mathcal{D}_{V}\!\!:\,\mathbb{R}^{2}\mapsto\mathbb{R}^{2}$. The neural map ${\mathcal{E}}$ takes in 640 real numbers associated with the modified phase at each frequency grid point, and it returns 3 real numbers associated with the 2 means of the latent space and the 1 standard deviation. The neural map $\mathcal{D}_{U}$ and $\mathcal{D}_{V}$, then, take 2 real numbers in the 2-D latent space and outputs two real numbers to be used as amplitudes and exponents in the pseudo-PN representation of the shape function. Each of these neural maps contains 4 fully connected hidden layers, and each hidden layer contains 512 neurons taking the ReLU activation [^4] Here, ReLU stands for the rectified linear unit, whose functional form is ${\rm ReLU}(x)=x\,\Theta(x)$ where $\Theta(\cdot)$ is the Heaviside step function. Other example activation functions include the hyperbolic tangent function, the sigmoid function, etc. The ReLU is the standard choice for deep learning because it mimics biological behavior in the cortex [102], while also addressing an issue common to neural networks, where numerical precision issues cause gradients to “vanish,” preventing further optimization [101]. Therefore, the entire encoder-decoder network contains 3,486,727 parameters that we must train by minimizing the loss function.